What is the number of square feet of steel plate required for a triangle with a height of 48" and a base of 66"
how do i figure this out???

Answer by Theo(13342)

the area of a triangle is equal to 1/2 * b * h.
b is the base
h is the height.
with a height of 48 inches and a base of 66 inches, the area of this triangle would be equal to 1/2 * 48 * 66 = 24 * 66 = 1584 square inches.